OpenStudy (anonymous):

Find a and d for the function f(x) =asin x +d such that the graph of f(x) matches the graph below.

OpenStudy (noelgreco):

The graph below?

OpenStudy (anonymous):

Sorry, I had to leave...I think I have the answer.

OpenStudy (anonymous):

I drew it before and had to redraw it...my laptop doesn't like this and keeps messing it up. I am switching computers. To do this correctly. But, here is what I came up with some guidance. f(x) asin x + d f(0) =1 a asin(0) + d =1

OpenStudy (anonymous):

more work to follow.

OpenStudy (anonymous):

cool much easier

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find a and d for the function f(x)= asin x + d such that the graph of f(x) matches the graph.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

f(0) =1 a sin (0) +d =1

OpenStudy (anonymous):

a (0) + d = 1 d=1

OpenStudy (anonymous):

a sin pi/2 + 1 =-1 a sin pi/2 =-2 -------- --- sin pi/2 sin pi/2 (on the left side sin pi/2) cancel a = -2/1=-2 a=-2 f(x) =-2sin x+1

OpenStudy (anonymous):

I guess I need a new tutor, lol.

OpenStudy (anonymous):

I just realized it would be good for me to know why you don't divide by one to find an unkown? If you get the time and feel like telling me...that would be excellent!